# Benchmarking Deep Learning-Based Methods for Irradiance Nowcasting with Sky Images

### v1

@@ 5.2 Ablation Experiments
Based on the results from the previous section, we then perform several ablation experiments using the ResNet50 model with the goal of investigating the impact that different configurations of pre-processing steps and training setups have on the performance. These configurations are summarized in Table 5. Note that, for the “$\Delta\text{t}$” and “Sample Interval” configuration, the searched values are only applied to the training set, while the default values are used for the test set.
@@ 5.2.1 Image Timestamps in $\mathcal{D}^{\text{folsom}}$
Table 6: Image timestamp alignment results for the four possible train/test combinations and the different test subsets in $\mathcal{D}^{\text{folsom}}$.
[…]
All File Name File Name 62.52 ± 0.18
File Name Date Modif. 56.72 ± 0.13
Date Modif. File Name 59.37 ± 0.34
Date Modif. Date Modif. 37.21 ± 0.28
Cloudy File Name File Name 89.76 ± 0.35
File Name Date Modif. 81.19 ± 0.18
Date Modif. File Name 85.34 ± 0.48
Date Modif. Date Modif. 52.62 ± 0.39
Clear File Name File Name 16.05 ± 0.60
File Name Date Modif. 15.76 ± 0.57
Date Modif. File Name 14.69 ± 0.43
Date Modif. Date Modif. 12.89 ± 0.43
As mentioned in subsection 3.6, aligning the sky images with the irradiance data in $\mathcal{D}^{\text{folsom}}$ using the image file names led to large label errors and the date modified of the images seemed to be more aligned with the irradiance data timestamps. To the best of our knowledge, this has not yet been reported in the literature so, in order to ascertain that the date modified is truly the correct timestamp for labeling the images in $\mathcal{D}^{\text{folsom}}$, a more quantitative analysis is also needed. To this end, we train two different models, one trained on a file name aligned training set and another on a date modified aligned training set. Both of these models are then evaluated on file named aligned testing set and a date modified aligned testing set, totaling four different train/test image timestamp alignment combinations. All other configurations are kept at their default values, depicted in Table 5.
As shown in Table 6, when both the training set and the testing set are aligned with the date modified timestamp the RMSE drops significantly, confirming the assumptions made in subsection 3.6. Perhaps unsurprisingly, this drop was much larger for the cloudy subset than the clear subset, since, for the majority of clear sky samples, a few minutes difference in alignment results in small GHI differences. On the other hand, under cloudy sky conditions, the sky image and irradiance alignment is much more critical.
Moreover, Table 6 also highlights the fact that evaluating with the date modified alignment is always better than with the file name alignment, even when training with the file name. The reason for this can be attributed to Figure 6, which shows that the file name alignment is much worse in 2016 than in 2014/2015. Since we selected 2014/2015 for training and 2016 for testing, this means that the model can still learn somewhat useful features if trained with the file name alignment. However, evaluation on the file name aligned test set will always result in very large errors, since the alignment is worst in 2016.
@@ 5.2.2 Time Shifting Operation
As described in subsection 3.5, we shift the irradiance measurements of the training set by applying a timestamp offset of $\Delta\text{t}$. For each dataset, we optimize $\Delta\text{t}\in\{-30\text{s}:10\text{s}:30\text{s}\}$ by applying a 5-fold CV (as described in subsection 4.1) and minimizing the validation loss. We then apply the optimal $\Delta\text{t}$ to the full training set to train the model and evaluate the results on the testing set.
The CV results are depicted in Figure 10, which shows the mean validation set RMSE across all 5 folds, for each dataset. Interestingly, in $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{nrel}}$, the validation RMSE decreases slightly for a negative $\Delta\text{t}$ and increases for a positive $\Delta\text{t}$, but in $\mathcal{D}^{\text{sirta}}$ the opposite happens. This suggests that the type of average applied to the GHI samples in $\mathcal{D}^{\text{folsom}}$ was the same as in $\mathcal{D}^{\text{nrel}}$, but different in $\mathcal{D}^{\text{sirta}}$. The optimal $\Delta\text{t}$ was -20s, 10s and -20s for $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively.
Regarding the results of the test set, Table 7 shows that using the optimal $\Delta\text{t}$ to train the models always leads to a consistent gain in performance for all datasets, especially in cloudy sky conditions, since the sky image and irradiance alignment is more critical during these periods.
@@ 5.2.3 Target Variable and Loss Function
NREL GHI 66.28 ± 0.68 82.10 ± 0.86 16.57 ± 0.19 36.05 ± 0.34 50.01 ± 0.54 11.57 ± 0.15
$k_{t}$ 62.23 ± 0.21 77.58 ± 0.26 10.46 ± 0.14 31.18 ± 0.22 44.66 ± 0.30 7.55 ± 0.14
$K_{t}$ 62.19 ± 0.32 77.44 ± 0.39 11.54 ± 0.10 31.36 ± 0.25 44.42 ± 0.37 8.46 ± 0.10
$k_{t}$ (weighted) 61.66 ± 0.10 76.87 ± 0.15 10.26 ± 0.59 31.02 ± 0.19 44.49 ± 0.12 7.40 ± 0.42
$K_{t}$ (weighted) 61.96 ± 0.32 77.18 ± 0.40 11.32 ± 0.16 31.34 ± 0.12 $4$ 0.13 8.36 ± 0.13
In this experiment, we consider five combinations of target variables and loss functions. Specifically, for the MSE loss (11), we consider three choices for the target variable $y\in\{I,k_{t},K_{t}\}$, resulting in the following expressions for the GHI estimate
@@
Although using the raw GHI value as target is the most natural and simple choice, it suffers from the fact that the GHI values will have different scales across different weather seasons. The left panel of Figure 11 depicts a $\Delta$GHI of almost 500 W/m² between the peak GHI value in summer and in winter, which, from their corresponding images, can be inferred based on the distance from the Sun center to the image center. This means that models trained to predict raw GHI values need to encode the Sun’s position in the image with extreme precision in order to be able to generalize it’s predictions, which can be a difficult task in cloudy sky conditions. By comparison, the $k_{t}$ value does not show this limitation (right panel of Figure 11), since the seasonal variance of the GHI values is already taken into account by the clear sky model. Another value that is not affected by this seasonal variance is the clearness index $K_{t}$, which just normalizes the GHI values by $I^{\text{extr}}$ instead of $I^{\text{clr}}$.
While the use of $k_{t}$ or $K_{t}$ as target solves the seasonal variance problem, it fails to take into account the diurnal variance of the GHI, i.e. a sample at the end of the day (with low GHI magnitude) will be weighted the same as a sample in the middle of the day (with high GHI magnitude). This is undesirable, since the middle of the day is the period with the higher load, hence samples during this period should be weighted more heavily. To solve this, we rescale the model prediction back into the GHI estimate before computing the squared error, leading to the loss function
@@
Figure 11: GHI and $k_{t}$ values comparison for a clear sky day showing the seasonal variance in $\mathcal{D}^{\text{folsom}}$. The left panel shows the GHI values and the right panel shows the $k_{t}$ values. The corresponding images to each labeled marker are shown in the middle panel.
The results for this experiment are summarized in Table 8, from which it is clear that modeling the raw GHI values yields the worst results by a large margin. This is especially true for $\mathcal{D}^{\text{folsom}}$, where there is almost a 50% improvement when using other target variables. Regarding the other target variables, the results are a little more inconclusive, since they are all within one standard deviation of each other and different targets being only slightly better in different datasets.
@@ 5.2.4 Camera Exposure
As shown in Figure 2, the images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ provide two different camera exposure levels, a long exposure form and a short exposure form. Although we use the long exposure form as our default value for all other experiments, the short exposure form provides more detail in the Sun disk area at the cost of providing less detail everywhere else. We hypothesize that in the nowcasting task, predicting the irradiance value given a short exposure image is a much easier task, since the only region of interest is the area around the Sun. To test this hypothesis, we train and evaluate models using the short exposure images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ and compare the results to their long exposure counterparts. $\mathcal{D}^{\text{folsom}}$ was not used in this experiment since it only provides one exposure form.
From the results presented in Table 9, using short exposure images increases the performance by a significant margin under cloudy sky conditions in $\mathcal{D}^{\text{sirta}}$, while maintaining almost the same performance under clear sky conditions. Since the majority of samples in this dataset are cloudy samples (see Table 2), the performance for the entire test set is also significantly better using short exposure images. However, in $\mathcal{D}^{\text{nrel}}$, it is the clear sky performance that increases, while the cloudy sky remains the same.
@@
NREL Long 62.10 ± 0.37 77.31 ± 0.50 11.63 ± 0.82 31.47 ± 0.32 44.54 ± 0.34 8.54 ± 0.72
Short 62.16 ± 0.20 77.56 ± 0.26 9.54 ± 0.22 30.43 ± 0.16 43.86 ± 0.24 6.88 ± 0.16
@@
NREL No Mask 66.35 ± 0.72 62.03 ± 0.33 36.28 ± 0.42 31.08 ± 0.25
Mask 65.57 ± 0.70 64.65 ± 0.40 35.52 ± 0.29 32.32 ± 0.12
@@ 5.2.5 Sample Interval
In this experiment we change the size of the training set by increasing the sample interval from 1 to 10 minutes. Since in $\mathcal{D}^{\text{nrel}}$ the minimum sample interval is 10 minutes, we do not search for any sample interval values using this dataset. All other configurations are kept at their defaults values. Our main goal with this experiment is to investigate if the larger error metrics in $\mathcal{D}^{\text{nrel}}$ are due to the higher sample interval in this dataset.
The results depicted in Figure 12 show that the test RMSE steadily increases with the larger sample intervals for the “All” and “Cloudy” subsets, however at a 10 minute interval the results for both $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{sirta}}$ are still far better than for $\mathcal{D}^{\text{nrel}}$, which indicates that the latter is a much harder dataset under cloudy sky conditions. However, for the “Clear” subset, a increasing the sample interval also deteriorates the performance, but much less so than in the other subsets. This is because clear sky conditions are very easy to model, such that little information will be gained by decreasing the sample interval. Interestingly, the test RMSE in $\mathcal{D}^{\text{nrel}}$ under clear sky conditions was very similar to the other two datasets.
@@ 5.2.6 Sun Mask
In Table 10 we show the nowcasting results for the different mask/target combinations, and compare it to [22]. Even though they used a different testing year for $\mathcal{D}^{\text{folsom}}$, we were also able to confirm, for all three datasets, that appending a Sun mask as a fourth channel to the input improves the model performance when the GHI is its target. However, when the $k_{t}$ is the target, the Sun mask actually degrades the model performance. We hypothesize that this is because of the fact that the Sun position in the image is less relevant when modeling the $k_{t}$, as discussed in subsubsection 5.2.3, such that appending a Sun mask would mostly add redundant information to the model, which could increase the risk of over-fitting the training set. Another possible explanation is the necessary redesign of the model to accept a 4-channel input, which forced a random weight initialization in the added weights of the first layer of the model. Since all other weights are initialized with their pre-trained values, this can disrupt the pre-learned low-level features extracted by the first layer.

### v2

@@ 5.2 Ablation Experiments
Based on the results from the previous section, we then perform several ablation experiments using the ResNet50 model with the goal of investigating the impact that different configurations of pre-processing steps and training setups have on the performance. These configurations are summarized in Table 6. Note that, for the $\Delta\text{t}$” and “Sample Interval” configuration, the searched values are only applied to the training set, while the default values are used for the test set.
@@ 5.2.1 Image Timestamps in $\mathcal{D}^{\text{folsom}}$
Table 7: GHI and DNI RMSE results for the four possible train/test combinations (FN = File Name; DM = Date Modified) and the different test subsets in $\mathcal{D}^{\text{folsom}}$.
RMSE $\text{W/m}^{2}$]
Sky Cond. Train Test GHI DNI
All FN FN 62.52 ± 0.18 96.04 ± 0.75
FN DM 56.72 ± 0.13 81.31 ± 1.28
DM FN 59.37 ± 0.34 96.60 ± 0.45
DM DM 37.21 ± 0.26 55.42 ± 0.94
Cloudy FN FN 89.76 ± 0.35 135.81 ± 0.94
FN DM 81.19 ± 0.18 113.88 ± 1.66
DM FN 85.34 ± 0.48 135.39 ± 0.36
DM DM 52.62 ± 0.35 73.25 ± 0.86
Clear FN FN 16.05 ± 0.60 33.13 ± 3.50
FN DM 15.76 ± 0.57 31.77 ± 3.49
DM FN 14.69 ± 0.43 37.55 ± 1.60
DM DM 12.89 ± 0.39 32.42 ± 1.67
As mentioned in subsection 3.6, aligning the sky images with the irradiance data in $\mathcal{D}^{\text{folsom}}$ using the image file names led to large label errors, while the date modified seemed much more aligned with the irradiance data. To the best of our knowledge, this has not yet been reported in the literature so, in order to ascertain that the date modified is truly the correct timestamp for labeling the images in $\mathcal{D}^{\text{folsom}}$, a more quantitative analysis is also needed. To this end, we train two different models, one trained on a file name aligned training set and another on a date modified aligned training set. Both of these models are then evaluated on file named aligned testing set and a date modified aligned testing set, totaling four different train/test image timestamp alignment combinations. All other configurations are kept at their default values, depicted in Table 6, and, for this particular experiment, we also evaluate the results on the DNI component, both to rule out the overirradiance hypothesis raised in subsection 3.6, as well as to offer a benchmark for future studies that use $\mathcal{D}^{\text{folsom}}$ to focus on this specific component.
Table 7 depicts the GHI and DNI results. One can see that, when the training set and the testing set are aligned with the date modified timestamp, the RMSE drops significantly, confirming the assumptions made in subsection 3.6. Perhaps unsurprisingly, this drop was much larger for the cloudy subset than the clear subset, since, for the majority of clear sky samples, a few minutes difference in alignment results in small GHI/DNI differences. On the other hand, under cloudy sky conditions, the sky image and irradiance alignment is much more critical.
Moreover, Table 7 also highlights the fact that evaluating with the date modified alignment is always better than with the file name alignment, even when training with the file name. The reason for this can be attributed to Figure 7, which shows that the file name alignment is much worse in 2016 than in 2014/2015. Since we selected 2014/2015 for training and 2016 for testing, this means that the model can still learn somewhat useful features when trained with the file name alignment.
@@ 5.2.2 Time Shifting Operation
As described in subsection 3.5, we shift the irradiance measurements of the training set by applying a timestamp offset of $\Delta\text{t}$. For each dataset, we optimize $\Delta\text{t}\in\{-30\text{s}:10\text{s}:30\text{s}\}$ by applying a 5-fold CV (as described in subsubsection 4.1.1) and minimizing the validation loss. We then apply the optimal $\Delta\text{t}$ to the full training set to train the model and evaluate the results on the testing set.
The CV results are depicted in Figure 13, which shows the mean validation set RMSE across all 5 folds, for each dataset. Interestingly, in $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{nrel}}$, the validation RMSE decreases slightly for a negative $\Delta\text{t}$ and increases for a positive $\Delta\text{t}$, but in $\mathcal{D}^{\text{sirta}}$ the opposite happens. This suggests that the type of average applied to the GHI samples in $\mathcal{D}^{\text{folsom}}$ was the same as in $\mathcal{D}^{\text{nrel}}$, but different in $\mathcal{D}^{\text{sirta}}$. The optimal $\Delta\text{t}$ was -20s, 10s and -30s for $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively.
Regarding the results of the test set, Table 8 shows that using the optimal $\Delta\text{t}$ to train the models always leads to a consistent gain in performance for all datasets, though in $\mathcal{D}^{\text{nrel}}$ this gain was close to negligible. The improvement is more apparent under cloudy sky conditions, since the sky image and irradiance alignment is more critical during these periods.
@@ 5.2.3 Target Variable and Loss Function
NREL GHI 42.43 ± 0.68 54.12 ± 0.90 14.04 ± 0.37 25.94 ± 0.36 36.64 ± 0.58 10.74 ± 0.20
$k_{t}$ 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
$K_{t}$ 38.20 ± 0.24 48.91 ± 0.36 11.60 ± 0.48 21.97 ± 0.11 31.16 ± 0.19 8.91 ± 0.40
$k_{t}$ (weighted) 38.06 ± 0.33 48.94 ± 0.43 10.18 ± 0.51 21.45 ± 0.30 31.11 ± 0.41 7.72 ± 0.25
$K_{t}$ (weighted) 38.04 ± 0.26 48.78 ± 0.35 11.13 ± 0.18 21.83 ± 0.28 31.15 ± 0.36 8.60 ± 0.19
In this experiment, we consider five combinations of target variables and loss functions. Specifically, for the MSE loss (12), we consider three choices for the target variable $y\in\{I,k_{t},K_{t}\}$, resulting in the following expressions for the GHI estimate:
@@
[…]
While the use of $k_{t}$ or $K_{t}$ as target solves the seasonal variance problem, it fails to take into account the diurnal variance of the GHI, i.e. a sample at the end of the day (with low GHI magnitude) will be weighted the same as a sample in the middle of the day (with high GHI magnitude). This could hurt the overall performance of the model, as it reduces the emphasis on middle-of-the-day samples, which are typically associated with larger prediction errors (RMSE and MAE). To solve this, we rescale the model prediction back into the GHI estimate before computing the squared error, leading to the loss function:
@@
Figure 14: GHI and $k_{t}$ values comparison for a clear sky day showing the seasonal variance in $\mathcal{D}^{\text{folsom}}$. The upper panel shows the GHI values and the lower panel shows the $k_{t}$ values. The corresponding images to each labeled marker are displayed on the right.
The results for this experiment are summarized in Table 9, from which it is clear that modeling the raw GHI values yields the worst results by a large margin. This is especially true for $\mathcal{D}^{\text{folsom}}$, where there is almost a 50% improvement when using other target variables. However, regarding the other target variables, the performance was largely the same, even when using the weighted loss function, which suggests that weighting midday samples more heavily may not be as beneficial as was initially hypothesized. Thus, we opt to use $k_{t}$ as our default target variable.
We also compare model performance across seasons and hours of the day, using either GHI or $k_{t}$ as target variable, and depict the results in Figure 15 and Figure 16. Since both the season and hour of day affect the magnitude of the GHI values, scale-dependent metric comparisons (such as RMSE and MAE) become more difficult. Consequently, we make use of the nRMSE metric for this analysis, due to its scale-invariant nature.
From Figure 15 one can see that using $k_{t}$ over GHI as the target variable yields a consistent gain in performance across all seasons in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, while in $\mathcal{D}^{\text{folsom}}$ this gain was slightly larger in spring and winter. We also note that the model performance tended to be best during summer and worse in winter. Since the summer season typically contains the highest number of clear sky samples, GHI estimation becomes much easier during this season. Additionally, extended daylight hours during summer increase the number of available sky images, which may introduce a slight model bias towards this season.
Regarding the hourly results depicted in Figure 16, we observe that models trained with $k_{t}$ target consistently outperform those trained with GHI throughout the day, with particularly notable improvements around sunrise and sunset. This is significant, as these periods typically coincide with large net load ramps, where accurate irradiance predictions are critical for efficient grid management.
Figure 15: Seasonal nRMSE for the GHI and $k_{t}$ target variables.
Figure 16: Hourly nRMSE for the GHI and $k_{t}$ target variables. For each hour mark, the nRMSE is computed across all samples between zero and 59 minutes.
@@ 5.2.4 Camera Exposure
As shown in Figure 2, the images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ provide two different camera exposure levels, a long exposure form and a short exposure form. Although we use the long exposure form as our default value for all other experiments, the short exposure form provides more detail in the Sun disk area at the cost of providing less detail everywhere else. We hypothesize that in the GHI estimation task, predicting the irradiance value given a short exposure image is a much easier task, since the main region of interest is the area around the Sun. To test this hypothesis, we train and evaluate models using the short exposure images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ and compare the results to their long exposure counterparts.
From the results presented in Table 10, using the short exposure images improved the performance on both datasets, especially under cloudy sky conditions. This behavior is expected, since the long exposure images can have saturated pixels around that Sun that make it hard to discern between a Sun and a cloud pixel. Although this confirms our hypothesis that the short exposure images yield better results on the GHI estimation task, we have not tested it on the forecasting task, which is why we default to the long exposure images in our other experiments.
@@
NREL Long 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
Short 37.34 ± 0.45 48.03 ± 0.61 9.86 ± 0.30 21.01 ± 0.22 30.52 ± 0.35 7.52 ± 0.32
@@
NREL No Mask 42.43 ± 0.68 38.75 ± 0.29 25.94 ± 0.36 21.73 ± 0.23
Mask 42.98 ± 0.52 41.62 ± 0.73 25.48 ± 0.40 22.83 ± 0.34
@@ 5.2.5 Sample Interval
In this experiment we change the size of the training set by increasing the sample interval from 1 to 10 minutes. Since in $\mathcal{D}^{\text{nrel}}$ the minimum sample interval is 10 minutes, we do not search for any sample interval values using this dataset. All other configurations are kept at their default values.
The results depicted in Figure 17 show that the test error rapidly increases with larger sample intervals under cloudy sky conditions, whereas under clear sky conditions the increase is much smaller. This behavior is expected since modeling clear sky conditions is a relatively simple task, simple enough that little new information would be gained by providing the model with additional clear sky training examples. In contrast, cloudy sky conditions can exhibit high variability due to complex cloud dynamics, shapes and overirradiance effects. As a result, decreasing the number of training examples rapidly decreases the model’s robustness under these conditions.
Additionally, it is also interesting to note that, when all datasets are on “equal ground” with a 10 minute sample interval, the behavior mentioned in subsection 5.1 is even more noticeable, that is, the “Cloudy” test set performance in $\mathcal{D}^{\text{folsom}}$ gets significantly worse by comparison. This further validates our hypothesis that the model is mistaking the lens contamination stains for false clouds.
@@ 5.2.6 Sun Mask
In Table 11 we show the GHI estimation results for the different mask/target combinations, and compare it to [22]. Even though they used a different testing year for $\mathcal{D}^{\text{folsom}}$, we were also able to confirm, for $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{sirta}}$, that appending a Sun mask as a fourth channel to the input improves the model performance when the GHI is its target, while doing so in $\mathcal{D}^{\text{nrel}}$ seemed to yield negligible results. However, when the $k_{t}$ is the target, the Sun mask actually degrades the model performance in all datasets. We hypothesize that this is because of the fact that the Sun position in the image is less relevant when modeling the $k_{t}$, as discussed in subsubsection 5.2.3, such that appending a Sun mask would mostly add redundant information to the model, which could increase the risk of over-fitting the training set. Another possible explanation is the necessary redesign of the model to accept a 4-channel input, which forced a random weight initialization in the added weights of the first layer of the model. Since all other weights are initialized with their pre-trained values, this can disrupt the pre-learned low-level features extracted by the first layer.
Figure 18: Ground truth and GHI predictions from the two-step and single-step approaches on two distinct future time windows: one where the Sun transitions from covered to visible (top panel), and the other where it transitions from visible to covered (bottom panel). In each panel, the left section displays the ASI context used as input for the forecasting model. The right section displays: ground truth future ASI (top row); ground truth and predicted future GHI values (middle row); predicted future ASI used as input for the GHI estimation model in the two-step approach (bottom row).
Table 12: Forecasting results for all models across all lead times. The proposed two-step approach showed significant improvements over the single-step approach, as well as the state of the art ECLIPSE [40] model. A lower bound on the forecasting error of the two-step approach is obtained by estimating the GHI using the ground truth instead of the predicted sky images.
RMSE [W/m2] (Forecast Skill [%])
2 min 4 min 6 min 8 min 10 min
Smart Persistence 93.6 (0.0%) 117.4 (0.0%) 129.8 (0.0%) 137.6 (0.0%) 143.1 (0.0%)
ECLIPSE [40] 83.8 (10.2%) - 98.5 (23.6%) - 109.1 (24.0%)
Single-step approach 90.5 (3.3%) 102.2 (13.0%) 108.2 (16.7%) 111.1 (19.3%) 113.9 (20.4%)
Two-step approach (Proposed) 80.9 (13.6%) 88.8 (24.3%) 94.6 (27.2%) 99.8 (27.5%) 103.9 (27.4%)
GHI estimation from real future ASI 44.1 (52.9%) 44.1 (62.4%) 44.1 (66.0%) 44.1 (67.9%) 44.1 (69.2%)
5.3 Forecasting Results
We begin our forecasting analysis by comparing the predictions from the two-step and single-step approaches on two distinct time windows. The first (top panel of Figure 18), where the Sun transitions from covered to visible, corresponds to the window with the highest standard deviation in ground truth future GHI values. The second (bottom panel of Figure 18), where the Sun transitions from visible to covered, corresponds to the window with the highest standard deviation in predicted future GHI values. Both cases show that the single-step approach tends to avoid large errors by simply predicting a mean GHI value, which comes at the cost of failing to predict GHI ramp events. In contrast, the two-step approach is capable of identifying such events, albeit with some limitations. By decoupling future event prediction from GHI estimation, the model is able to capture the cloud dynamics more effectively, enabling it to correctly predict when the Sun will be covered and when it will be visible. Consequently, when the generated images are combined with a GHI estimation model, the two-step approach yields much more accurate and sensible predictions.
However, we note that the generated images are far from realistic, exhibiting noticeable blur and heavy artifacts along the patch borders. The blur is likely due to our choice of loss function, which was only the pixel-wise MSE, while the border artifacts likely stems from our choice of linear decoder, which independently decodes each patch embedding. Nevertheless, we believe these results provide a strong proof of concept for the two-step approach in GHI nowcasting.
In Table 12 we compare the test set results for all forecasting models. These results further corroborate our proposed two-step approach, since it shows a significant improvement over the other models across all lead times. However, the lower bound forecasting error, which we obtain by replacing the predicted sky images with the ground truth before the GHI estimation module, indicates that there is still a large performance gap to be gained by improving the forecasting module and generating more realistic images.